# Jet-shaped filamentary ejecta in common envelope evolution

### v1

@@ 1 Introduction
In a common envelope evolution (CEE), a compact object, the secondary, orbits inside the extended envelope of the primary star. The interaction of the secondary star with the envelope is not symmetrical, i.e., it has no cylindrical symmetry because of the density gradient in the envelope and the circular, elliptical, or spiraling-in orbit. Consequently, the mass that the secondary accretes has a net-specific angular momentum. If the secondary is sufficiently compact, the accretion process proceeds through an accretion disk around the secondary. As some theoretical studies (e.g., Armitage & Livio 2000; Chevalier 2012; López-Cámara et al. 2019, 2020) showed, such accretion disks in CEE might launch jets, mainly when the secondary is a neutron star (NS) or a black hole (for a recent review see Grichener 2024). Also, main sequence stars in CEE might also launch jets (e.g., Blackman & Lucchini 2014; Soker 2023). The jets remove mass from the secondary vicinity, hence reducing mass accretion rate, and by that, operate in a negative feedback mechanism, e.g., Soker (2016b) for an old review, and Grichener, Cohen, & Soker (2021) and Hillel, Schreier, & Soker (2022) for one-dimensional (1D) and 3D simulations, respectively, of this negative feedback mechanism. Alongside the negative, there is a positive feedback mechanism because the jets that the accretion disk launches carry energy that reduces the pressure around the secondary (e.g., Shiber, Schreier, & Soker 2016; Chamandy et al. 2018; López-Cámara et al. 2019, 2020).
@@
Most 3D simulations of the CEE do not include jets that the secondary star launches (e.g., Passy et al. 2012; Ricker & Taam 2012; Nandez et al. 2014; Staff et al. 2016a; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017b; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier & Pejcha 2024; González-Bolívar et al. 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024, for a small fraction of papers presenting CEE simulation; Roepke & De Marco 2023 for a review). The small number of 3D hydrodynamical studies of the CEE and the grazing envelope evolution that includes the jets that the secondary launches have to give up other ingredients of the CEE (e.g., Moreno Méndez et al. 2017; Shiber 2018; Shiber & Soker 2018; López-Cámara et al. 2019; Schreier et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Lopez-Camara et al. 2022; Zou et al. 2022; Soker 2022; Hillel, Schreier, & Soker 2023; Schreier, Hillel, & Soker 2023; Gurjar et al. 2024; Shiber & Iaconi 2024). Unrelated to our present study are simulations that obtain collimated outflows from the distorted envelope at the final phases of the CEE (e.g., Zou et al. 2020; Moreno et al. 2022; Ondratschek et al. 2022; Vetter et al. 2024), as suggested many years ago by an analytical calculation (Soker, 1992).
@@ 2 The numerical setup
The first ingredient is the non-rotating stellar model. We use a zero-age-main-sequence star of metalicity $Z=0.02$ and mass $M_{\rm 1,ZAMS}=15M_{\odot}$ evolved to the RSG phase using the one-dimensional (1D) stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019). Our primary focus is 3D hydrodynamical simulations using flash (Fryxell et al., 2000). We transport the 1D stellar model to the 3D numerical grid when its mass and radius are $M_{1}=12.5M_{\odot}$ and $R_{\rm RSG}=881\,R_{\odot}$, respectively.
The inner $20\%$ of the stellar radius is kept fixed. We use a stationary gravitational field computed at the initial time from the stellar model, including the fixed inner sphere. Our numerical grid is a Cartesian grid with outflow boundary conditions. The matter occupying the grid is an ideal gas ($\gamma=5/3$), including radiation pressure.
@@
The third ingredient is the launching of jets. The orbiting NS does not affect the simulation other than these jets launched from its location at any moment. Our scheme does not change the mass density of any computational cell but changes the velocity and internal energy of the existing material near the NS. The power of the two opposing jets varies with the location of the NS according to the Bondi-Hoyle-Lyttleton (BHL) mass accretion rate from the unperturbed envelope reduced by a multiplication factor $\zeta\simeq 0.002-0.005$ (Grichener, Cohen, & Soker 2021; Hillel, Schreier, & Soker 2022). In this study, we take $\zeta=7.5\times 10^{-5}$ because too strong jets dominate and obscure the effect of the envelope’s rotation. This energy and momentum are inserted in a jet-envelope interaction region, which is assumed to be a cylinder. The scheme conserves mass, momentum, and energy.
@@
where $M$ and $R$ are the mass and stellar radius at the beginning of the simulation. In addition to the non-rotating case (Hillel, Schreier, & Soker, 2023), we simulate two cases with $\alpha=0.25$ and $\alpha=0.5$.
@@ 3 The rotating three-dimensional stellar model
We describe results using regular and high resolutions. In the regular resolution simulations, we apply adaptive mesh refinement with the smallest cell size in the simulations being $L_{\rm G}/128=3.90625\times 10^{12}{~{}\rm cm}$, where $L_{G}$ is the size of the simulation box. In the high-resolution simulations, all grid cells have a size of $L_{\rm G}/256=1.95\times 10^{12}{~{}\rm cm}$. If we do not indicate the resolution, it is the regular one. Because of limited computational resources, we use only the regular resolution when we include jets (Section 4). We take the $z$ axis along the initial envelope angular momentum.
@@
In our study of the non-rotating 3D model (Hillel, Schreier, & Soker, 2023), we found that the envelope performs two oscillations, then decays. We argued that this is an expected behavior of RSGs, and therefore, there is no need to stabilize the model before starting the CEE simulations; just let the star oscillate. We follow the initial mass in spherical shells to reveal the oscillatory behavior. We mark each initial spherical shell with a ‘tracer;’ all cells in a shell start at $t=0$ with ${\rm tracer}=1$. The tracer decreases as the initial shell’s material mixes with gas outside the shell. The tracer is a fraction of each cell’s mass originating in the shell. The tracer is always between zero and one. The two initial shells we follow are $600R_{\odot}<r<650R_{\odot}$ and $800R_{\odot}<r<850R_{\odot}$; the stellar radius when we start the 3D simulation is $R_{\rm RSG}=881\,R_{\odot}$.
Figure 1 presents, by dashed lines in three colors, the average radius of the two shells for three simulations with different initial rotation velocities according to equation (1); all lines in Figure 1 are with regular resolution as they compare the simulations without jets with those with jets that we study in Section 4. We learn that the rotating models also perform two oscillations before decaying. As expected, the centrifugal force increases the average radius. This larger radius increases the dynamical time and the oscillation period, as we learn from the peaks in the oscillating dashed lines. We reiterate our claim that there is no need to relax the 3D stellar model; just let it oscillate. Introducing an energy source in the inner 3D stellar model that mimics the nuclear burning in the core might postpone the oscillation decay. This is the topic of ongoing research. Figure 2 compares the high with the regular resolution simulation, both without jets. The differences are relatively small, showing that it is adequate to use the regular resolution when we add jets (as we have insufficient computer resources to simulate with the high resolution when we include jets).
Figure 1: The average radii of the tracers of two initial spherical shells $800R_{\odot}<r<850R_{\odot}$ (upper figure) and $600R_{\odot}<r<650R_{\odot}$ (lower figure). The solid lines are for the simulations with jets, and the dashed lines are for those without jets. The insets show the initial envelope rotation according to equation (1). The Keplerian orbital period on the surface of the unperturbed RSG is $P_{\rm Kep}=2.35{~{}\rm yr}$, and the dynamical time is $P_{\rm D}\equiv(G\bar{\rho})^{-1/2}=0.76{~{}\rm yr}$ where $\bar{\rho}$ is the average density of the initial RSG model.
@@
Figure 2: The average radii of the tracers of two initial spherical shells $800R_{\odot}<r<850R_{\odot}$ and $600R_{\odot}<r<650R_{\odot}$ similar to Figure 1, but comparing the regular resolution (thin-dashed lines) and the high resolution (thick-solid lines) of no-jet simulations of a rotating envelope with $\alpha=0.5$. The differences are sufficiently small to justify simulating jets with the regular resolution.
@@
The upper four panels of Figure 4 present the tracer maps in the meridional plane $x=0$ of the two shells that we show in Figure 1 for the $\alpha=0.5$ simulation, and at four times as in Figure 3, including minima and maxima of the average outer shell radius. The upper left panel presents the two shells at the beginning of the simulation ($t=12{~{}\rm days}$). The two shells are well separated with envelope gas free of the shells’ material. At $t=0.92{~{}\rm yr}$, the first maximum in the oscillation, the envelope flattens, as expected from the centrifugal force, and the shells’ material mixes with the region between the shells. This mixing is expected as the envelope is convective, as demonstrated in the non-rotating case in the previous paper (Hillel, Schreier, & Soker, 2023). While the inner shell maintains its identity, the outer shell smears to larger radii than its initial radius. The left panel in the second row at $t=1.8{~{}\rm yr}$ shows that at the minimum of the oscillation, some material from the outer shell spreads to the edge of the grid while some outer shell material rebuilds the shell near the equatorial plane. The inner shell continues to mix with its surroundings (color change from deep red at $t=0.92{~{}\rm yr}$ to yellow or faint red) but still maintains its identity. The right panel in the second row shows the continuation of this evolution. The two panels in the third row of Figure 4 show the inner shell at $t=2.8{~{}\rm yr}$ (second maximum) and at $t=3.5{~{}\rm yr}$, respectively. We can see that throughout the entire simulation, the inner shell remains confined. The lower two panels present tracer maps of the high-resolution simulation at $t=2.7{~{}\rm yr}$: lower left: the tracer of the outer shell; lower right: the tracer of the inner shell. The shells’ behavior shows that although the outer part expands and its material flows out of the grid boundary, the inner parts stay inside the star. This behavior implies we can trust the simulations if the companion spirals into the envelope. First, the energy the jets deposit significantly affects mass ejection and envelope disruption more than oscillation and numerical inaccuracies. Second, in a short time, the NS spirals to the center, where the model stays intact for several years, which is several dynamical times, more than we need for our present study with jets (Section 4).
@@
Figure 4: Tracer maps of two shells in the meridional $x=0$ plane at $t=12{~{}\rm days}$ (top left), $t=0.92{~{}\rm yr}$ (first maximum in the oscillator motion; top right panel), $t=1.8{~{}\rm yr}$ (first minimum; second left), and $t=2.7{~{}\rm yr}$ (second right). The two panels in the third raw present only the tracer of the inner shell at $t=2.7{~{}\rm yr}$ (left), and $t=3.5{~{}\rm yr}$ (right), at about the dynamical time. The upper six panels are of the regular simulation with $\alpha=0.5$. The last two panels show the tracers of the outer (left) and inner (right) shells at $t=2.7{~{}\rm yr}$ of the high-resolution simulation with $\alpha=0.5$. Units on the axes are cm. The tracer scales are on top of the panels and change between panels.
@@
Figure 5 presents the velocity maps in a quarter of the meridional plane $x=0$ and at two times, for the regular (upper panels) and high (lower panel) resolution no-jet simulations. The blue zone is of no interest to us as it is the low-density medium that we insert around the star at $t=0$ (and the arrows are hard to see). This figure emphasizes the non-radial oscillation. At $t=0.92{~{}\rm yr}$ (left panels), the material along the polar direction near $z=60\times 10^{12}{~{}\rm cm}$ moves inward (contracts), while material near the equatorial plane ($y\simeq 110\times 10^{12}{~{}\rm cm}$ in the maps) expands. At $t=2.3{~{}\rm yr}$ (right panels) the velocities in these regions are in the opposite directions. There are large vortices that are clearer in the high-resolution simulation. These four maps also show that the differences between the regular and high-resolution simulations are not large.
@@
Figure 5: Velocity vectors on top of density maps in a quarter of the meridional $x=0$ plane at $t=0.92{~{}\rm yr}$ (left) and at $t=2.3{~{}\rm yr}$ (right). The upper and lower panels are for the regular and high-resolution simulations for $\alpha=0.5$. Units on the axes are cm, and the largest arrow at $t=0.92{~{}\rm yr}$ and at $t=2.3{~{}\rm yr}$ corresponds to velocities of $90{~{}\rm km}{~{}\rm s}^{-1}$ and $50{~{}\rm km}{~{}\rm s}^{-1}$, respectively. The color bars depict the density from $10^{-12}{~{}\rm g}{~{}\rm cm}^{-3}$ (deep blue) to $10^{-8}{~{}\rm g}{~{}\rm cm}^{-3}$ (deep red).
@@
We can summarize the simulations of the stellar model before the onset of the CEE by strengthening our claim from Hillel, Schreier, & Soker (2023) that there is no need to stabilize and relax the stellar model before starting the CEE simulation. The very large oscillations and the mixing in the envelope that results from the turbulent convection are actual processes. RSG stars have strong envelope convection and prominent amplitude pulsations. The radius variation (maximum radius minus minimum radius in an oscillation cycle) of observed long-period variable stars can be of the order of the stellar radius $\Delta R\simeq R$ (e.g., Trabucchi et al. 2021). Our code is inaccurate in calculating the very outer parts of the stellar model after about a dynamical time $P_{\rm D}\equiv(G\bar{\rho})^{-1/2}=0.76{~{}\rm yr}$. However, if we start the CEE simulation immediately when we start the simulation, without relaxation time, the effect of the companion is much more significant than the inaccuracies of our numerical code.

### v2

@@ 1 Introduction
In a common envelope evolution (CEE), a compact object, the secondary, orbits inside the extended envelope of the primary star. The interaction of the secondary star with the envelope is not symmetrical, i.e., it has no cylindrical symmetry because of the density gradient in the envelope and the circular, elliptical, or spiraling-in orbit. Consequently, the mass that the secondary accretes has a net-specific angular momentum. If the secondary is sufficiently compact, the accretion process proceeds through an accretion disk around the secondary. As some theoretical studies (e.g., Armitage & Livio 2000; Chevalier 2012; López-Cámara et al. 2019, 2020) showed, such accretion disks in CEE might launch jets, mainly when the secondary is a neutron star (NS) or a black hole (for a recent review see Grichener 2024). Also, main sequence stars in CEE might also launch jets (e.g., Blackman & Lucchini 2014; Soker 2023). The disk around the NS might continuously exist from the pre-CEE phase, like from the Roche lobe overflow process (e.g., Juarez-Garcia et al. 2025). The jets remove mass from the secondary vicinity, hence reducing mass accretion rate, and by that, operate in a negative feedback mechanism, e.g., Soker (2016b) for an old review, and Grichener, Cohen, & Soker (2021) and Hillel, Schreier, & Soker (2022) for one-dimensional (1D) and 3D simulations, respectively, of this negative feedback mechanism. Alongside the negative, there is a positive feedback mechanism because the jets that the accretion disk launches carry energy that reduces the pressure around the secondary (e.g., Shiber, Schreier, & Soker 2016; Chamandy et al. 2018; López-Cámara et al. 2019, 2020).
@@
Most 3D simulations of the CEE do not include jets that the secondary star launches (e.g., Passy et al. 2012; Ricker & Taam 2012; Nandez et al. 2014; Staff et al. 2016a; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017b; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier & Pejcha 2024; González-Bolívar et al. 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024, 2025, for a small fraction of papers presenting CEE simulation; Roepke & De Marco 2023 for a review). The small number of 3D hydrodynamical studies of the CEE and the grazing envelope evolution that includes the jets that the secondary launches have to give up other ingredients of the CEE (e.g., Moreno Méndez et al. 2017; Shiber 2018; Shiber & Soker 2018; López-Cámara et al. 2019; Schreier et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Lopez-Camara et al. 2022; Zou et al. 2022; Soker 2022; Hillel, Schreier, & Soker 2023; Schreier, Hillel, & Soker 2023; Gurjar et al. 2024; Shiber & Iaconi 2024). Unrelated to our present study are simulations that obtain collimated outflows from the distorted envelope at the final phases of the CEE (e.g., Zou et al. 2020; Moreno et al. 2022; Ondratschek et al. 2022; Vetter et al. 2024), as suggested many years ago by an analytical calculation (Soker, 1992).
@@ 2 The numerical setup
The first ingredient is the non-rotating stellar model. We use a zero-age-main-sequence star of metalicity $Z=0.02$ and mass $M_{\rm 1,ZAMS}=15M_{\odot}$ evolved to the RSG phase using the one-dimensional (1D) stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019). Our primary focus is 3D hydrodynamical simulations using flash (Fryxell et al., 2000). We transport the 1D stellar model to the 3D numerical grid when its mass and radius are $M_{1}=12.5M_{\odot}$ and $R_{\rm RSG}=881\,R_{\odot}$, respectively. Cells outside the stellar model are initialized with a low density of $\rho_{\rm grid,0}=2.1\times 10^{-13}{~{}\rm g}{~{}\rm cm}^{-3}$ and a temperature of $T_{\rm grid,0}=1100{~{}\rm K}$, and outflow boundary conditions are used.
The inner $20\%$ of the stellar radius is kept fixed. This inner inert core, with a radius of $R_{\rm inert}=176R_{\odot}$, saves large amounts of computational resources, as simulating inner envelope zones requires smaller time steps. We use a stationary gravitational field computed at the initial time from the stellar model, including the fixed inner sphere. Our numerical grid is a Cartesian grid with outflow boundary conditions. The matter occupying the grid is an ideal gas ($\gamma=5/3$), including radiation pressure.
@@
The third ingredient is the launching of jets. The orbiting NS does not affect the simulation, except for the jets launched from its location at any given moment. Our scheme does not alter the mass density of any computational cell, but rather changes the velocity and internal energy of the existing material near the NS. The power of the two opposing jets varies with the location of the NS according to the Bondi-Hoyle-Lyttleton (BHL) mass accretion rate from the unperturbed envelope reduced by a multiplication factor $\zeta\simeq 0.002-0.005$ (Grichener, Cohen, & Soker 2021; Hillel, Schreier, & Soker 2022). The mass and radius of the NS are taken to be $M_{\rm NS}=1.4M_{\odot}$ and $R_{\rm NS}=12{~{}\rm km}$, respectively. In this study, we take $\zeta=7.5\times 10^{-5}$ for three reasons. (1) Much stronger jets will dominate and obscure the effect of the envelope’s rotation. (2) The topic of this study is the effect of envelope rotation, so we simulate three cases with different rotation velocities. We do not scan other parameters, such as elliptical orbits and jets’ power (3). As we found in an earlier paper (Hillel, Schreier, & Soker, 2022), simulations with much more powerful jets are highly demanding and will take months to simulate. Future studies should simulate other parameters for specific needs as observations of other numerical studies appear and include radiative transport, a significant process for CEE of massive stars (e.g., Lau et al. 2025).
The energy and momentum of the jets are inserted in a jet-envelope interaction region, which is assumed to be a cylinder with a base radius of $4\times 10^{12}{~{}\rm cm}$ and total height $14\times 10^{12}{~{}\rm cm}$ The scheme conserves mass, momentum, and energy.
@@
where $M$ and $R$ are the mass and stellar radius at the beginning of the simulation. In addition to the non-rotating case (Hillel, Schreier, & Soker, 2023), we simulate two cases with $\alpha=0.25$ and $\alpha=0.5$. For $\alpha=0.5$, on the equator, the centrifugal force is one-quarter of the gravitational force. We did not simulate cases with faster rotation because, in such cases, centrifugal forces become too large to justify the assumption that the star is spherical, despite its rotation. Simulating a much faster rotation requires starting the simulation earlier and having a higher resolution near the equatorial plane. We do not yet have the required numerical resources.
@@ 3 The rotating three-dimensional stellar model
We present results obtained using both regular and high resolutions. the center of the grid is at the center of the RSG stellar model. In the regular resolution simulations, we use a uniform grid with a cell size of $L_{\rm G}/128=3.90625\times 10^{12}{~{}\rm cm}$, where $L_{G}=5\times 10^{14}$ is the size of the simulation cube . In the high-resolution simulations, the cells have a size of $L_{\rm G}/256=1.95\times 10^{12}{~{}\rm cm}$. If we do not specify the resolution, it defaults to the regular one. Due to limited computational resources, we use only the regular resolution when including jets (Section 4). We take the $z$-axis along the initial envelope’s angular momentum.
@@
In our study of the non-rotating 3D model (Hillel, Schreier, & Soker, 2023), we found that the envelope performs two oscillations and then decays. We argued that this is an expected behavior of RSGs, and therefore, there is no need to stabilize the model before starting the CEE simulations; just let the star oscillate. We follow the initial mass in spherical shells to reveal the oscillatory behavior. We mark each initial spherical shell with a ‘tracer’; all cells in a shell start at $t=0$ with ${\rm tracer}=1$. The tracer decreases as the initial shell’s material mixes with gas outside the shell. The tracer is a fraction of each cell’s mass originating in the shell. The tracer value is always between 0 and 1. The two initial shells we follow are $600R_{\odot}<r<650R_{\odot}$ and $800R_{\odot}<r<850R_{\odot}$; the stellar radius when we start the 3D simulation is $R_{\rm RSG}=881\,R_{\odot}$.
Figure 1 compares the high-resolution simulation with the regular-resolution simulation, both without jets. This figure shows the average radii of the tracers of two initial spherical shells $800R_{\odot}<r<850R_{\odot}$ and $600R_{\odot}<r<650R_{\odot}$. The oscillation of the envelope for two cycles before decaying is evident. The period between the two maxima is $\simeq 1.7{~{}\rm yr}$. For comparison, the Keplerian orbital period on the surface of the unperturbed RSG is $P_{\rm Kep}=2.35{~{}\rm yr}$, and the dynamical time is $P_{\rm D}\equiv(G\bar{\rho})^{-1/2}=0.76{~{}\rm yr}$ where $\bar{\rho}$ is the average density of the initial RSG model. The oscillations are of a dynamic nature (see below). The differences between the regular and high resolutions are relatively small. The Amplitude from the first minimum at $t\simeq 1.7{~{}\rm yr}$ to the second maximum at $t\simeq 2.6{~{}\rm yr}$ of the regular resolution is 1.22 times larger than that of the high resolution. The average radius in the high-resolution is smaller. The ratio of this amplitude to the average radius at the minimum and second maximum in the regular resolution is 1.16 larger than that of the high resolution. The oscillation period is larger in the regular resolution as the star expands more than in the high resolution; therefore, the dynamical time is larger. These small differences, $\simeq 20\%$, indicate that it is sufficient to use the regular resolution when we add jets (as we have insufficient computer resources to simulate with the high resolution when we include jets).
@@
Figure 1: The average radii of the tracers of two initial spherical shells $800R_{\odot}<r<850R_{\odot}$ and $600R_{\odot}<r<650R_{\odot}$, comparing the regular resolution (thin-dashed lines) and the high resolution (thick-solid lines) of no-jet simulations of a rotating envelope with $\alpha=0.5$. The differences are sufficiently small to justify simulating jets with the regular resolution.
Figure 2 presents, by dashed lines in three colors, the average radius of the two shells for three simulations with different initial rotation velocities according to equation (1); all lines in Figure 2 are with regular resolution as they compare the simulations without jets with those with jets that we study in Section 4. We learn that the rotating models also perform two oscillations before decaying. As expected, the centrifugal force increases the average radius. The larger radius increases the dynamical time and the oscillation period, as indicated by the peaks in the oscillating dashed lines. We reiterate our claim that there is no need to relax the 3D stellar model; it can simply be allowed to oscillate. Introducing an energy source in the inner 3D stellar model that mimics the nuclear burning in the core might postpone the oscillation decay. This is the topic of ongoing research.
Figure 2: The average radii of the tracers for two initial spherical shells: $800R_{\odot}<r<850R_{\odot}$ (upper figure) and $600R_{\odot}<r<650R_{\odot}$ (lower figure). The solid lines represent simulations with jets, and the dashed lines represent those without jets. The insets show the initial envelope rotation according to equation (1). The Keplerian orbital period on the surface of the unperturbed RSG is $P_{\rm Kep}=2.35{~{}\rm yr}$, and the dynamical time is $P_{\rm D}\equiv(G\bar{\rho})^{-1/2}=0.76{~{}\rm yr}$ where $\bar{\rho}$ is the average density of the initial RSG model.
The solid lines in Figure 2 show that the jets do not prevent stellar oscillations. There are oscillations convolved with envelope expansion. The deposition of energy by the jets causes a significant expansion of the envelope, which increases the oscillation period. The oscillations are prominent in the outer tracer shell (upper panel of Figure 2). The first maximum is at a time similar to that of the simulations without jets, but the radius is larger. The second minimum of the oscillation appears as a very shallow increase in radius. The second peak is clear in the $\alpha=0$ and $\alpha=0.25$ cases. Still, it occurs much later compared to the jetless simulations, as the envelope is large and the dynamic time is significantly longer.
Let us elaborate on the nature of the oscillations in the 3D stellar model and their relation to those of RSG stars. The driving energy of the oscillations in giant stars is the stellar luminosity resulting from nuclear burning in the core, such as the $\kappa$ mechanism, where opacity increases during compression and decreases during expansion. We do not yet have an energy source near the center of our 3D stellar model to mimic the stellar luminosity. Therefore, the oscillations we find here cannot be the same as those of RSG stars. The oscillations of the 3D stellar model result from the deviations of envelope zones from hydrostatic equilibrium. Indeed, these oscillations decay after about two periods. What we claim is that the oscillations resulting from the departure of the envelope from exact hydrostatic equilibrium mimic, in their periods and amplitudes, real RSG stellar oscillations more closely than a fully relaxed model does. We obtain prominent radial oscillations. The flow of the envelope mass in the non-rotating stellar models of Hillel, Schreier, & Soker (2023) also exhibits tangential motions. The tangential motions are mainly the result of the large convective vortices as the envelope is also convective. Whether there are also non-radial pulsations unrelated to the convection is difficult to tell from our analysis and requires a separate study that better resolves the convective motion. Had we relaxed the model before starting the simulation, the stellar model would not oscillate, as evident from the decay of the oscillations. To have oscillations even in a relaxed model requires the numerical code to include an energy source near the center that mimics nuclear energy in the core, an envelope model with opacity that depends on the local temperature and density, and following ionization fraction.
@@
The upper four panels of Figure 4 present the tracer maps in the meridional plane $x=0$ of the two shells that we show in Figure 2 for the $\alpha=0.5$ simulation, and at four times as in Figure 3, including minima and maxima of the average outer shell radius. The upper left panel presents the two shells at the beginning of the simulation ($t=12{~{}\rm days}$). The two shells are well separated with envelope gas free of the shells’ material. At $t=0.92{~{}\rm yr}$, the first maximum in the oscillation, the envelope flattens, as expected from the centrifugal force, and the shells’ material mixes with the region between the shells. This mixing is expected as the envelope is convective, as demonstrated in the non-rotating case in the previous paper (Hillel, Schreier, & Soker, 2023). As we simulate only the middle and outer zones of the RSG envelope, and not the inner inert sphere, at $t=0$, the entire envelope (beside a very thin layer near the surface) is convective. While the inner shell maintains its identity, the outer shell smears to larger radii than its initial radius. The left panel in the second row at $t=1.8{~{}\rm yr}$ shows that at the minimum of the oscillation, some material from the outer shell spreads to the edge of the grid while some outer shell material rebuilds the shell near the equatorial plane. The inner shell continues to mix with its surroundings (color change from deep red at $t=0.92{~{}\rm yr}$ to yellow or faint red) but still maintains its identity. The right panel in the second row shows the continuation of this evolution. The two panels in the third row of Figure 4 show the inner shell at $t=2.8{~{}\rm yr}$ (second maximum) and at $t=3.5{~{}\rm yr}$, respectively. We can see that throughout the entire simulation, the inner shell remains confined. The lower two panels present tracer maps of the high-resolution simulation at $t=2.7{~{}\rm yr}$: lower left: the tracer of the outer shell; lower right: the tracer of the inner shell. The shells’ behavior shows that although the outer part expands and its material flows out of the grid boundary, the inner parts stay inside the star. This behavior implies that we can trust the simulations if the companion spirals into the envelope. First, the energy the jets deposit significantly affects mass ejection and envelope disruption more than oscillation and numerical inaccuracies. Second, in a short time, the NS spirals to the center, where the model stays intact for several years, which is several dynamical times, more than we need for our present study with jets (Section 4).
@@
Figure 4: Tracer maps of two shells in the meridional $x=0$ plane at $t=12{~{}\rm days}$ (top left), $t=0.92{~{}\rm yr}$ (first maximum in the oscillator motion; top right panel), $t=1.8{~{}\rm yr}$ (first minimum; second left), and $t=2.7{~{}\rm yr}$ (second right). The two panels in the third row present only the tracer of the inner shell at $t=2.7{~{}\rm yr}$ (left), and $t=3.5{~{}\rm yr}$ (right), at about the dynamical time. The upper six panels are of the regular simulation with $\alpha=0.5$. The last two panels show the tracers of the outer (left) and inner (right) shells at $t=2.7{~{}\rm yr}$ of the high-resolution simulation with $\alpha=0.5$. The units on the axes are cm. The tracer scales are on top of the panels and change between panels.
@@
Figure 5 presents the velocity maps in a quarter of the meridional plane $x=0$ and at two times, for the regular (upper panels) and high (lower panel) resolution no-jet simulations. The blue zone is of no interest to us as it is the low-density medium that we insert around the star at $t=0$ (and the arrows are hard to see). This figure emphasizes the non-radial oscillation. At $t=0.92{~{}\rm yr}$ (left panels), the material along the polar direction near $z=60\times 10^{12}{~{}\rm cm}$ moves inward (contracts), while material near the equatorial plane ($y\simeq 110\times 10^{12}{~{}\rm cm}$ in the maps) expands. At $t=2.3{~{}\rm yr}$ (right panels), the velocities in these regions are in the opposite directions. There are large vortices that are clearer in the high-resolution simulation. These four maps also show that the differences between the regular and high-resolution simulations are not large.
@@
Figure 5: Velocity vectors on top of density maps in a quarter of the meridional $x=0$ plane at $t=0.92{~{}\rm yr}$ (left) and at $t=2.3{~{}\rm yr}$ (right). Note that we simulate the entire volume, but for clarity, present only one-quarter of the plane. The upper and lower panels are for the regular and high-resolution simulations for $\alpha=0.5$. Units on the axes are cm, and the largest arrow at $t=0.92{~{}\rm yr}$ and at $t=2.3{~{}\rm yr}$ corresponds to velocities of $90{~{}\rm km}{~{}\rm s}^{-1}$ and $50{~{}\rm km}{~{}\rm s}^{-1}$, respectively. The color bars depict the density from $10^{-12}{~{}\rm g}{~{}\rm cm}^{-3}$ (deep blue) to $10^{-8}{~{}\rm g}{~{}\rm cm}^{-3}$ (deep red).
@@
We can summarize the simulations of the stellar model before the onset of the CEE by strengthening our claim from Hillel, Schreier, & Soker (2023) that there is no need to stabilize and relax the stellar model before starting the CEE simulation. The very large oscillations and the mixing in the envelope that results from the turbulent convection are actual processes. RSG stars have strong envelope convection and prominent amplitude pulsations. The radius variation (maximum radius minus minimum radius in an oscillation cycle) of observed long-period variable stars can be of the order of the stellar radius $\Delta R\simeq R$ (e.g., Trabucchi et al. 2021). Our code is inaccurate in calculating the very outer parts of the stellar model after about a dynamical time $P_{\rm D}\equiv(G\bar{\rho})^{-1/2}=0.76{~{}\rm yr}$. The code does much better and for several dynamical times, which is about five years, in most of the envelope. Since we launch very energetic jets as we start the simulations, with no delay time, the effect of the jets in expelling the outer parts of the envelope is much larger than the numerical inaccuracy in the outer parts. In other words, the jets expel the outer parts of the envelope before the numerical inaccuracies develop.